# excitations, and scale dependence

### v1

@@ 3.3 Halo Abundance and Bias
We now check the impact of PNGs on statistics of massive halos. While PNGs impact halos over a range of mass scales, the resolution of our simulation limits us to studying only objects with $M_{\rm 200c}>10^{13.5}\,{\rm M}_{\odot}/h$ as these are resolved by at least 30 particles. This resolution is adequate for our target of measuring halo counts and halo clustering. The analysis in this section follows that of Anbajagane et al. (2023) and Paper I who identified the formation of massive halos as the primary origin for why the weak lensing field is sensitive to $f_{\rm NL}$, in agreement with earlier work on the PNG information content in halo statistics (Dalal et al. 2008; Shirasaki et al. 2012; Marian et al. 2011; Hilbert et al. 2012).
@@
Figure 6 presents the change in the halo mass function (HMF) for a unit change in $f_{\rm NL}$. As shown in Anbajagane et al. (2023), the sensitivity of the HMF to PNGs is an order of magnitude larger than the other effects discussed above. Following the arguments of prior work (e.g., Dalal et al. 2008; Shirasaki et al. 2012; Anbajagane et al. 2023, Paper I), we note that PNGs change the shape of density field’s 1-point distribution. This change preferentially up/down-weights the tails of the distribution, and the abundance of massive halos is sensitive to these tails (Press & Schechter 1974). As a result, the HMF changes significantly as we vary the amplitude of PNGs. Following the same argument, the most massive halos (whose formation depends on the more extreme/low-probability regime of the distribution) are more significantly impacted by a PNG template relative to their less-massive counterparts.
@@ 3.4 Interplay Between the Primordial Bispectrum and Power Spectrum
The constraints on $A_{\rm pk}$ are also interesting on their own, as a single-parameter analysis. Forecasts of linear resonances generally consider $w_{\rm pk}\gtrsim 10$ (Beutler et al. 2019b, see their Figure 10) due to modeling challenges. For example, the forecasted uncertainty from galaxy clustering probes increases by an order of magnitude below $\lesssim 30\,{\rm Mpc}$, due to the challenge of modeling the galaxy field on such scales using perturbation theory. According to Beutler et al. (2019b), Stage IV spectroscopic surveys will achieve $\sigma(A_{\rm pk})\sim 0.3$. When considering only lensing data, we find $w=8$ could be constrained at $\sigma(A_{\rm pk})=0.03$, which is an order of magnitude better. However, the situation is inverted for larger frequencies—our constraints for $w=64$ are $\sigma(A_{\rm pk})=0.05$, whereas the aforementioned surveys achieve $\sigma(A_{\rm pk})=0.001$ in this regime. Similar to Section 3.1 above, we find lensing is most powerful when constraining features that originate on small scales.
Appendix A provides a more detailed description of how power spectrum oscillations alter nonlinear structure formation. For intuition on the nature of our constraints in Figure 8, we now provide a simple example using linear theory. Figure 9 shows the resonance-modified signal in the linear matter power spectrum (left), the corresponding real-space correlation function (middle), and the standard deviation of the matter density fluctuations (right). The latter is connected to the formation of massive halos and is useful for understanding how these oscillations will impact halo abundances (Press & Schechter 1974). In all cases, we rescale the power spectrum (after the feature has been added) such that it produces $\sigma_{8}=0.834$. All calculations are performed using the Core Cosmology Library (Chisari et al. 2019).
@@
In summary, Figure 9 identifies why lensing data is more sensitive to $w_{\rm pk}\lesssim 30$ and why it is more sensitive to linear resonance models than logarithmic resonance ones. Finally, we note these results use linear theory (without any nonlinear extensions or simulations) and are only meant to illustrate the main features of the underlying signal. The more rigorous analysis of these signals, folding in the full nonlinear behavior via $N$-body simulations, is performed in Appendix A. Previous works have also studied different models of power spectrum oscillations using $N$-body simulations (e.g., Schaeffer & Schneider 2021; Stahl et al. 2025).
@@ 4 Conclusions
NN-body simulations have been an ever-present tool in the analysis of density fluctuations in the late Universe. Their uses in our fiducial cosmology analyses range from model testing and validation (e.g., DeRose et al. 2019; To et al. 2024), to semi-analytic extensions into more nonlinear scales (e.g., Takahashi et al. 2012, 2020), and to full forward-modeled simulation-based inference (e.g., Zürcher et al. 2022; Fluri et al. 2022; Gatti et al. 2024; Prat et al. 2025). Different aspects of these techniques have been used extensively to model probes like weak lensing, cluster abundance, galaxy clustering etc. (e.g., Krause et al. 2021; Anbajagane et al. 2025c; To et al. 2025; Gomes et al. 2025). However, such techniques have generally not been viable for studies of inflation—and particularly for studies beyond the simpler inflationary models—due to the lack of simulation suites that include this signal. We have improved on this in Paper I, by providing a framework for generating simulations with arbitrary bispectra in their initial conditions. In that work, we utilized this method to show the utility of late-time, nonlinear probes in studying the signals from particle interactions in the primordial Universe; the so-called “cosmological collider physics” models. In this work, we have shown the same techniques enable us to study a wider range of models, and highlight how nonlinear structure can be more than competitive in constraining the model parameter spaces.
@@ A.2 Halo Abundance and Bias
The HMF shows clear, non-monotonic features as a function of halo mass. Focusing on the linear resonance models, we see that increasing $w_{\rm pk}$ causes the derivative to be suppressed for less massive halos. In the logarithmic models, the derivatives are highly suppressed and essentially zero. Following Press & Schechter (1974), the HMF depends on the integral of the power spectra, i.e., the $\sigma(r)$ quantity presented in the right panels of Figure 9. Rapid oscillations in $P(k)$ get averaged out in this integral and result in little-to-no sensitivity of the HMF to the resonance signals.

### v2

@@ 3.3 Halo Abundance and Bias
We now check the impact of PNGs on statistics of massive halos. While PNGs impact halos over a range of mass scales, the resolution of our simulation limits us to studying only objects with $M_{\rm 200c}>10^{13.5}\,{\rm M}_{\odot}/h$ as these are resolved by at least 30 particles. This resolution is adequate for our target of measuring halo counts and halo clustering. The analysis in this section follows that of Anbajagane et al. (2023) and Paper I who identified the formation of massive halos as the primary origin for why the weak lensing field is sensitive to $f_{\rm NL}$, in agreement with earlier work on the PNG information content in halo statistics (Dalal et al. 2008; Shirasaki et al. 2012; Marian et al. 2011; Hilbert et al. 2012; Jung et al. 2023).
@@
Figure 6 presents the change in the halo mass function (HMF) for a unit change in $f_{\rm NL}$. As shown in Anbajagane et al. (2023), the sensitivity of the HMF to PNGs is an order of magnitude larger than the other effects discussed above. Following the arguments of prior work (e.g., Dalal et al. 2008; Shirasaki et al. 2012; Anbajagane et al. 2023, Paper I), we note that PNGs change the shape of density field’s 1-point distribution. This change preferentially up/down-weights the tails of the distribution, and the abundance of massive halos is sensitive to these tails (Press and Schechter 1974). As a result, the HMF changes significantly as we vary the amplitude of PNGs. Following the same argument, the most massive halos (whose formation depends on the more extreme/low-probability regime of the distribution) are more significantly impacted by a PNG template relative to their less-massive counterparts.
@@ 3.4 Interplay Between the Primordial Bispectrum and Power Spectrum
The constraints on $A_{\rm pk}$ are also interesting on their own, as a single-parameter analysis. Forecasts of linear resonances generally consider $w_{\rm pk}\gtrsim 10$ (Beutler et al. 2019a, see their Figure 10) due to modeling challenges. For example, the forecasted uncertainty from galaxy clustering probes increases by an order of magnitude below $\lesssim 30\,{\rm Mpc}$, due to the challenge of modeling the galaxy field on such scales using perturbation theory. According to Beutler et al. (2019a), Stage IV spectroscopic surveys will achieve $\sigma(A_{\rm pk})\sim 0.3$. When considering only lensing data, we find $w=8$ could be constrained at $\sigma(A_{\rm pk})=0.03$, which is an order of magnitude better. However, the situation is inverted for larger frequencies—our constraints for $w=64$ are $\sigma(A_{\rm pk})=0.05$, whereas the aforementioned surveys achieve $\sigma(A_{\rm pk})=0.001$ in this regime. Similar to Section 3.1 above, we find lensing is most powerful when constraining features that originate on small scales.
Appendix A provides a more detailed description of how power spectrum oscillations alter nonlinear structure formation. For intuition on the nature of our constraints in Figure 8, we now provide a simple example using linear theory. Figure 9 shows the resonance-modified signal in the linear matter power spectrum (left), the corresponding real-space correlation function (middle), and the standard deviation of the matter density fluctuations (right). The latter is connected to the formation of massive halos and is useful for understanding how these oscillations will impact halo abundances (Press and Schechter 1974). In all cases, we rescale the power spectrum (after the feature has been added) such that it produces $\sigma_{8}=0.834$. All calculations are performed using the Core Cosmology Library (Chisari et al. 2019).
@@
In summary, Figure 9 identifies why lensing data is more sensitive to $w_{\rm pk}\lesssim 30$ and why it is more sensitive to linear resonance models than logarithmic resonance ones. Finally, we note these results use linear theory (without any nonlinear extensions or simulations) and are only meant to illustrate the main features of the underlying signal. The more rigorous analysis of these signals, folding in the full nonlinear behavior via $N$-body simulations, is performed in Appendix A. Previous works have also studied different models of power spectrum oscillations using $N$-body simulations (e.g., Schaeffer and Schneider 2021; Stahl et al. 2025).
@@ 4 Conclusions
NN-body simulations have been an ever-present tool in the analysis of density fluctuations in the late Universe. Their uses in our fiducial cosmology analyses range from model testing and validation (e.g., DeRose et al. 2019; To et al. 2024), to semi-analytic extensions into more nonlinear scales (e.g., Takahashi et al. 2012, 2020), and to full forward-modeled simulation-based inference (e.g., Zürcher et al. 2022; Fluri et al. 2022; Gatti et al. 2024; Prat et al. 2025). Different aspects of these techniques have been used extensively to model probes like weak lensing, cluster abundance, galaxy clustering etc. (e.g., Krause et al. 2021; Anbajagane et al. 2025b; To et al. 2025; Gomes et al. 2025). However, such techniques have generally not been viable for studies of inflation—and particularly for studies beyond the simpler inflationary models—due to the lack of simulation suites that include this signal. We have improved on this in Paper I, by providing a framework for generating simulations with arbitrary bispectra in their initial conditions. In that work, we utilized this method to show the utility of late-time, nonlinear probes in studying the signals from particle interactions in the primordial Universe; the so-called “cosmological collider physics” models. In this work, we have shown the same techniques enable us to study a wider range of models, and highlight how nonlinear structure can be more than competitive in constraining the model parameter spaces.
@@ A.2 Halo Abundance and Bias
The HMF shows clear, non-monotonic features as a function of halo mass. Focusing on the linear resonance models, we see that increasing $w_{\rm pk}$ causes the derivative to be suppressed for less massive halos. In the logarithmic models, the derivatives are highly suppressed and essentially zero. Following Press and Schechter (1974), the HMF depends on the integral of the power spectra, i.e., the $\sigma(r)$ quantity presented in the right panels of Figure 9. Rapid oscillations in $P(k)$ get averaged out in this integral and result in little-to-no sensitivity of the HMF to the resonance signals.